OpenStudy (anonymous):

What is the solution to the equation x – 9 = –4? A. –5 B. –36 C. 5 D. –13

OpenStudy (the_fizicx99):

Add 9 to both sides

OpenStudy (the_fizicx99):

x - 9 = -4 +9 +9

OpenStudy (anonymous):

ok thanks

OpenStudy (the_fizicx99):

What did you get for x? Yw! :p

OpenStudy (anonymous):

a

OpenStudy (the_fizicx99):

How did you get A .. (-4) + 9 = ?

OpenStudy (the_fizicx99):

Lets use a number line, your on -4, add 4 to your -4 then you''l be at 0. You still have a remaining 5.. move those 5 to the right (positive). Which is?